# Biomechanics-Aware Trajectory Optimization for Navigation during Robotic Physiotherapy

### v1

@@ III-D Computational performance (active and passive subject)
We also used simulations to compare the computational efficiency of our choice to capture the human skeletal dynamics with OpenSimAD, against numerically querying the original OpenSim model during the optimization. This was done for an active subject as well as for a passive subject, modulating the time horizon for the trajectory optimization instances (4). In Tab. I, we report the average computation times for solving (4) when the system dynamics are enforced at the collocation points through the two different methods. The difference is roughly two orders of magnitude in favor of BATON, and real-time trajectory optimization was possible only with OpenSimAD, granting re-planning capabilities at a frequency of roughly $10$ Hz with appropriate selection of the planning horizon and its discretization. BATON achieved the same computational performances during the robot experiments.
@@
IV Discussion
Figure 9: Different supraspinatus activation $\alpha_{\text{ss}}$ causes different strain landscapes for the corresponding tendon. Left: no activation - right: activation is increased, causing the lower-strain path to be different and leading the human through different shoulder poses. The same value of shoulder elevation is used to compare the two trajectories, showing a difference of more than $10^{\circ}$ visualized both on the model and our human subject.
@@ IV Discussion
We have developed a new approach to enable robot-based physiotherapy called BATON, which can automate and personalize rehabilitation through the navigation of underlying biomechanical outcomes. Achieving our first research aim, BATON directly couples human musculoskeletal outcomes to a robot controller through trajectory optimization. Embedding high-fidelity musculoskeletal outcomes is crucial for enabling the robotic system to deliver rehabilitation for rotator cuff injuries, where insights into patient biomechanics are paramount. Knowledge of the patient’s range of motion and tissue loads is essential for optimizing safe movements that avoid high-strain configurations associated with high risks of re-injury. These biomechanical insights allowed us to plan and execute robotic rehabilitation trajectories, but can also be leveraged by experienced therapists to improve therapy’s efficacy and safety.
To accomplish our second aim of real-time replanning capabilities, we decoupled the navigation of musculoskeletal loading into tissue loading landscapes as strain maps and algorithmically differentiable human dynamics for control. The tissue-loading landscape represents the navigable rehabilitation space, like a map of underwater reefs used to help a ship safely navigate, and the skeletal dynamics of the human shoulder describe the motion response of the vessel. This unique approach allows us to take advantage of the details of high-fidelity biomechanical models without facing prohibitive computational costs during real-time control. In both passive and active scenarios, we found BATON to operate several orders of magnitudes faster than using traditional musculoskeletal modeling and simulation techniques (Tab. I). As a result, real-time re-planning of reference trajectories for the rehabilitation robot was attained at an average update rate of just below 10 Hz. This allowed us to close the loop between sensing, optimization, and reference tracking both in simulation, where multiple edge-cases were investigated (Fig. 7 and Fig. 8), as well as on the real robot (Fig. 9). Besides improved computational efficiency, decoupling also provides a more modular overview of different aspects of the musculoskeletal system. For example, strain maps can provide an intuitive visual interpretation of strain mechanics as an indicator of re-injury risk during shoulder therapy.
Further, we analyzed the performances of BATON through simulations and physical experiments and achieved the third aim of this study. The optimized trajectories are tunable to achieve a trade-off between the minimization of tendon strain and the directness of the path (Fig. 4). We guaranteed a high-quality reproduction of the optimal path executed by the robot, including effective support of the subject’s arm (Fig.5 and Fig.6). Through a combination of simulation and physical experiments, we demonstrated how BATON can quickly adapt to a changing strain landscape due to unforeseen changes in rotator cuff muscle activity.
Overall, our results demonstrate how embedding a high-fidelity representation of the human musculoskeletal system in the controller improves physical human-robot interaction in the context of automated yet responsive rehabilitation. This unlocks the design of robotic controllers that can monitor musculoskeletal quantities that even experienced therapists lack access to, potentially enhancing the utility of these systems in clinical rehabilitation. The next step is to evaluate the system on target user and patient groups.
@@
The different modalities (i.e., an active and a passive subject) explored in this study can provide solutions for different stages of the therapy. After surgery and in the initial phases of treatment after an injury, the patient is generally restricted in their mobility, and movements are carefully applied by the therapist who supports the arm and shoulder. We were able to control robot-applied movements over long-horizon trajectories to increase the range of motion of the shoulder while minimizing tendon strain in the rotator cuff to avoid re-injury (Fig. 5). As treatment progresses and as the patient regains strength and mobility, patients become and are encouraged to be more active during the exercises, in which case navigation must be more responsive. We tested rapid changes in muscle activation simulating patient reactions, to show that high-strain regions (even due to emerging activation) are successfully avoided using the active real-time re-planning of our biomechanics-aware trajectory optimization methods.
@@
While this study demonstrated and evaluated key aspects of the novel method, we acknowledge that there are some limitations that require interesting future work. We described the state of the human shoulder focusing on the mobility of the glenohumeral joint alone, and planned human motions along two DoFs ($PE$ and $SE$). In our setting, the motion along the third DoF ($AR$) of the human glenohumeral joint was not optimized, and the current value of $AR$ determined the strain map on which the optimization took place. This simplification allowed us to validate the core concept, but spatial (3D) rehabilitation trajectories should be investigated for real-world applications. For example, while the strain of the rotator cuff tendon depends on the relative position between the upper arm and the shoulder blade, other movement DoFs, such as the scapula mobility or the human torso orientation, can also influence tissue loading in the shoulder. The human pose estimation that we employed did not include other DoFs and may rely on simplifying assumptions that are too restrictive for a wide range of physical therapy applications. While an accurate estimation of the subject’s body motion was beyond the scope of this work, better methods to monitor human posture and movement could be integrated. Viable options could rely on motion capture techniques, which offer high precision but are difficult to deploy, while video camera systems or IMUs might be more practical but also less reliable. Apart from resulting in more complex computations, a challenging aspect of treating additional DoFs is delivering visual information that a patient/therapist can easily interpret. In the future, we will explore how to extend BATON to also include additional DoFs, such as those related to the scapula, with an approach to measure the state of the scapula itself.
Another limitation is related to estimating realistic muscle activations. We simulated muscle activations to test extreme conditions for BATON and verify that the planned trajectories remained in low-strain regions to give us confidence that the automated responses are safe. Additionally, using simulated activations ensured the repeatability and reproducibility of the responses. As such, while these activations may not be physiological, including these extremes allowed us to test the worst-case scenario of abrupt human responses and show BATON’s potential to be safe under these conditions. It is important to note that the rotator cuff muscles are not accessible by common surface electromyography, and direct measurement would be impossible during physiotherapy. In the future, we plan to integrate our novel rapid model-based techniques (e.g., RMR solver [22]) into BATON so that muscle activities can be estimated reliably. Since the RMR solver requires only human pose and interaction forces as input, it is a very promising candidate for estimating physiologically plausible muscle forces and activations. The formulation of the RMR solver would enable us to retain real-time capabilities and to account for activation dynamics as well. When considering real muscle activations, resorting to predictive methods like BATON for synthesizing safe physical human-robot interaction will be even more important, allowing a robot to plan movements to minimize other relevant physical quantities such as interaction forces between the human and the robot, and joint reaction forces.
V Conclusion
We have presented a novel approach to physical human-robot interaction that leverages biomechanical-aware trajectory optimization to generate and execute rehabilitation trajectories applied by a real robot to a human subject, leveraging a personalized musculoskeletal model. We demonstrated BATON for planning robot-assisted movements that account for the mechanical strain in the rotator cuff tendons, both during the early stages of rehabilitation when the human is passively guided, and as muscle activation emerges when the subject actively responds during in therapy. In contrast to previous robotic approaches where computationally expensive biomechanical metrics were not considered, we took the first step in unlocking the use of high-fidelity musculoskeletal models in real-time biomechanics-aware robotic controllers. While our work focuses on physiotherapy robotics, the ability of automated systems to estimate and respond to the changing internal states of the human body can resonate far beyond physical therapy. Our approach can enable the development of the next generation of biomechanical-aware controllers, and pass the “baton” to researchers in related fields of ergonomics, exoskeleton control, movement disorders, etc., where automatic, responsive, and personalized robotics can transform outcomes.

### v2

@@ V-D Computational performances
Finally, we used simulations to compare the computational efficiency of our choice to capture the human skeletal dynamics with OpenSimAD, against numerically querying the original OpenSim model during the optimization (Sec. II-A). This was done for an active subject as well as for a passive subject, modulating the time horizon for the trajectory optimization instances (7). The average computation times for solving (7) when the system dynamics are enforced at the collocation points through the two different methods indicate a difference of roughly two orders of magnitude in favor of BATON (Tab. I). Real-time trajectory optimization was possible only with OpenSimAD, granting re-planning capabilities at a frequency of roughly $10$ Hz with appropriate selection of the planning horizon and its discretization. BATON achieved the same computational performances during the robot experiments.
@@
Figure 8: Different infraspinatus activation $\alpha_{\text{is}}$ causes different strain map topography for the corresponding tendon in our physical human-robot interaction experiments. Feeding human pose and interaction torque to BATON, it adapts to changes in estimated muscle activity. Here, we reveal the progression of rehabilitation trajectories along varying strain maps for our subject complying with the robot guidance (top row) and voluntarily activating their infraspinatus (second row). The solid blue lines represent the trajectory executed for each case, while dashed white lines allow us to compare them to the other case. Bottom rows: evolution of estimated infraspinatus activation $\hat{\alpha}_{\text{is}}$, and interaction torque modulated by the subject during the motion for both cases.
VI Discussion
VI-A Motion Planning on strain maps
@@
BATON is a new approach that enables robot-assisted physiotherapy to both automate and personalize rehabilitation by navigating underlying human biomechanical outcomes during therapy delivery. Achieving our first research aim (A1), BATON directly embeds human biomechanics into a robot controller through trajectory optimization and adapts to muscle activation changes as they occur. Accounting for high-fidelity musculoskeletal outcomes is crucial to deliver effective robotic rehabilitation. We developed and applied BATON to address the complexities and challenges of traditional physical therapy for rotator cuff injuries, where insights into patient-level tendon strains are paramount. Knowledge of the patient’s tissue loads and range of motion is essential for optimizing safe movements that avoid high-strain configurations associated with risks of re-injury. These biomechanical insights allow us to plan and execute robotic rehabilitation trajectories, but can also be harnessed by physical therapists to guide the efficacy and safety of applied movements.
In both passive and active scenarios, we found BATON to operate several orders of magnitude faster than using traditional musculoskeletal modeling and simulation techniques (Tab. I). As a result, real-time re-planning of reference trajectories for the rehabilitation robot was attained at an average update rate of just below 10 Hz.
VI-B Dynamic adaptation to the human actions
An integral part of the safety and potential autonomy of BATON is the feedback of the human response into the planning and control loop. We demonstrate the estimation and adaptation to changes in muscle activity in both simulation (Figs 6-7) and activity inducing movements in a test subject (Fig. 8). The online model-based estimates of muscle activity account for changes in the patient’s unpredictable volitional actions and/or reflexive responses due to physical human-robot interaction, achieving our second aim (A2). As a consequence, our system can inform the planner about underlying muscle activity, e.g., rotator cuff muscles, which cannot be measured by surface EMG. Although the proposed approach does not require surface EMG, such data could be used, if available, to augment the solver in estimating muscle activity.
We analyzed our algorithm over multiple edge cases in simulation (Fig. 6 and Fig. 7), to verify the capability of our online trajectory optimization to cope with extreme strain map variations (potentially due to muscle activation changes). Then, we leveraged a model-based muscle estimation algorithm to close the loop between sensing, online trajectory optimization, and reference tracking on the real robot (Fig. 8). The use of the RMR solver enabled our system to estimate physiological muscle recruitment without the use of instrumenting our subject with any kind of sensor (e.g., surface EMG ones), decreasing the time required for setup and enabling simultaneous monitoring of virtually any muscle in the human body (even deep stabilizers like the rotator cuff). Considering the infraspinatus muscle as a target tissue for our real-robot experiments, we achieved online therapy adaptations to its activation changes during physical human-robot interaction (Sec. V-C). Muscle activation was monitored both when the subject complied with the robot motion commands, and when they voluntarily elicited muscle activation. In the first case, rather small variations occurred, and they were likely due to the mechanical role of the infraspinatus as a glenohumeral stabilizer during humeral elevation. In the second case, voluntary muscle activation was estimated, in a scenario similar to real therapy when a patient’s neuromuscular reflexes might be triggered, or the human might interact in unexpected ways with the robot.
VI-C Evaluation insights of BATON performance
Further, we analyzed the performances of BATON through simulations and physical experiments and achieved the third aim of this study (A3). The optimized trajectories can be tuned to reach a trade-off between the minimization of tendon strain and the directness of the path (Fig. 4). BATON’s optimized trajectories guaranteed lower instantaneous and cumulative tendon strain than the SoTA planner [38], which instead traversed higher strain regions (Fig. 5). Moreover, BATON produced smoother trajectories, while the SoTA planner delivered trajectories presenting higher oscillations, potentially in contrast with natural human movement preferences [52]. These oscillations are intrinsic to the assumptions of the SoTA kinematic planner, as it operates over a discrete version of the strain maps, resulting in various non-differentiable points along the commanded reference. Beyond resulting in higher shoulder accelerations, the SoTA planner also applied higher peak forces to the human arm in our experiments (Fig. 5, bottom row), a result that can be attributable to ignoring human dynamics during planning.
@@ VI-C Evaluation insights of BATON performance
The different modalities (i.e., an active and a passive subject) explored in this study can provide solutions for different stages of the therapy. After surgery or in early treatment after an injury, the patient is generally restricted in their mobility, and movements are carefully applied by the therapist who supports the arm and shoulder. We were able to control robot-led movements over long-horizon trajectories to increase the range of motion of the shoulder while minimizing tendon strain in the rotator cuff to avoid re-injury (Fig. 5). As treatment progresses and as the patient regains strength and mobility, patients become and are encouraged to be more active during the exercises, in which case navigation must be more responsive (Figs. 6-7-8).
In conclusion, our results demonstrate that embedding a high-fidelity representation of the human musculoskeletal system in the controller improves physical human-robot interaction in the context of automated yet responsive rehabilitation. In this study, we focused on rotator cuff strain monitoring and minimization, but other quantities could also be included (such as joint reaction forces or muscle activations themselves). Overall, the use of high-fidelity biomechanical models unlocks the design of robotic controllers that can monitor musculoskeletal quantities that even experienced therapists lack access to, potentially enhancing the utility of these systems in clinical rehabilitation. The next step is to evaluate the system on target users and patient groups.
VI-D Limitations and future work
@@
We demonstrated and evaluated key aspects of BATON that provide confidence in its application for physiotherapy in the future. Nonetheless, we acknowledge that there are several limitations that require further investigation. First, we focused on the mobility of the glenohumeral joint alone, and planned human motions along two of its DoFs ($PE$ and $SE$). In our setting, the motion along the third DoF ($AR$) of the human glenohumeral joint was not optimized, and the current value of $AR$ determined the strain map on which the optimization took place. This simplification allowed us to validate the core concept, but spatial (3D) rehabilitation trajectories should be investigated for use in therapeutic applications. For example, the strain of the rotator cuff tendon depends on the relative position between the upper arm and the shoulder blade, and other movement DoFs, such as the scapula mobility or the human torso orientation, will also influence tissue loading in the shoulder.
Second, the human pose estimation that we employed did not include the complete human state and may rely on simplifying assumptions that are too restrictive for a wide range of physical therapy applications. While an accurate estimation of the subject’s body motion was beyond the scope of this work, better methods to monitor human posture and movement could be integrated. Viable options could rely on motion capture techniques, which offer high precision but are difficult to deploy, while video camera systems or IMUs might be more practical but also less reliable. Exoskeletons could also provide better human kinematics with respect to the collaborative robotics arm employed in this study. In the future, we plan to explore how to extend BATON to also include additional DoFs, such as those related to the scapula, with an approach to measure the state of the scapula itself. However, apart from resulting in more complex computations, a challenging aspect of treating additional DoFs is delivering visual information that a patient/therapist can easily interpret.
A third limitation lies in the strain value delivered by the shoulder model we selected. While the model has been validated in previous studies for what concerns the accuracy of shoulder kinematics [53] and muscle activity [40, 27], a thorough validation of the estimated tendon behavior across the model’s range of motion is necessary before we can apply robotic physical therapy for rehabilitation from rotator cuff tendon tears on a patient population.
Lastly, we did not predict human behavior at planning time, opting instead for online adaptation to volitional actions and reflexive reactions. While BATON could be extended to include predictive models of human motion or neuromuscular dynamics(e.g., [54, 55]), they would provide only uncertain predictions that deviate from actual human behavior. The fundamental challenge of reliably predicting human decisions and behaviors remains. As such, we designed BATON to use actual behavior, necessary to ensure safety and effective performance, prioritizing real-time responsiveness to changes in human muscle activations estimated from measured movement and interaction forces. Future advances in biomechanical estimation could further enhance BATON’s performance.
VII Conclusion
We presented BATON, a novel approach to physical human-robot interaction and its application for shoulder rotator cuff physiotherapy. Through biomechanics-aware trajectory optimization, BATON generates and executes rehabilitation trajectories applied to a human subject by a collaborative robot, leveraging a personalized musculoskeletal model. Mechanical tendon strain was monitored and minimized both during the early, passive stages of rehabilitation, and as muscle activation emerges when the subject responds, either actively or reactively, during therapy.
Unlike prior work in robotic-assisted physiotherapy, which does not directly monitor internal biomechanical metrics due to modeling and computational complexities, we have taken the first step to unlock the power of high-fidelity musculoskeletal models for real-time biomechanics-aware robotic controllers. Moreover, integral to BATON is an online alternative to sensor-based muscle activity estimations, such as surface EMG, which cannot measure underlying muscles like the rotator cuff.
While our primary focus is robotic rehabilitation, the ability of automated systems to estimate and respond to the changing internal states of the human body resonates far beyond physical therapy. BATON represents a foundational shift toward the next generation of human-centered assistive and augmentative robotics. We pass the “baton” to researchers and engineers in ergonomics, exoskeleton control, movement disorders, and beyond, where awareness of human biomechanics and subject/patient response is needed for robots to transform outcomes.
Appendix
Experimentally, commanding the Cartesian end-effector pose through (15) produces non-negligible tracking errors along the vertical axis if the human subject cannot support the weight of their own arm. Since full-arm support is typical of early rotator cuff rehabilitation [56], we implemented a gravity-compensation strategy modifying (15) to include the generalized torques $\boldsymbol{u}^{*}$ into ${}^{\text{base}}\bar{\boldsymbol{p}}_{\text{EE}}$. Let us define the end-effector Cartesian position as ${}^{\text{base}}\bar{\boldsymbol{p}}_{\text{EE}}={}^{\text{base}}\bar{\boldsymbol{p}}_{\text{GH}}+{}^{\text{base}}\bar{\boldsymbol{R}}_{\text{EE}}{}^{\text{base}}\bar{\boldsymbol{p}}_{\text{GH}}=[p_{x}\ p_{y}\ p_{z}]^{T}$, similar to (15). From our biomechanics-aware trajectory optimization (Sec. II-B) we obtain the human generalized torques needed to track such a reference at every time instant. Thus, to deliver the required model-based torque along the $SE$ DoF to the human subject, we added to the vertical EE reference a displacement $\delta_{z}$:
$\delta_{z}=\frac{u^{*,\text{SE}}}{K_{z}L_{\text{hum}}\mathrm{sin}(\hat{\text{SE}})}$ (17)
resulting in ${}^{\text{base}}\bar{\boldsymbol{p}}_{\text{EE}}=[p_{x},\ p_{y},\ p_{z}+\delta_{z}]^{T}$
This approach is personalizable to each subject through their specific biomechanics and appears more robust to model inaccuracies than purely torque-based compensation methods (i.e. [57]). We show experimentally that this adjustment achieves the desired vertical EE setpoint, fully supporting the human arm without resorting to impractically high stiffness in the controller.
Fig. 9 shows the Cartesian reference positions for the robot EE together with the actual ones across the duration of the experiment. In particular, the bottom row of the figure presents the tracking performance along the robot’s vertical axis. It can be noted that the vertical reference $z_{\text{}ref}=p_{z}+\delta_{z}$, accounting for the optimized offset in (17), is not reached. However, this allows the Cartesian impedance controller to provide the correct support, such that the optimal end-effector height $z_{\text{opt}}=p_{z}$ is tracked.
Figure 9: Tracking performances of BATON’s optimal trajectory during passive therapy, expressed into the Cartesian coordinates of the robotic end-effector ${}^{\text{base}}\boldsymbol{p}_{\text{EE}}$. The bottom graph highlights the effect of using the personalized musculoskeletal model to inform the gravity compensation of the human arm: the offset reference given to the controller (solid black) makes it so that the current position (dashed) better tracks the optimal one (solid blue).